In an article appearing in Today’s Health, a writer states that a single serving of popcorn has on average 75 calories. To determine if the average number of calories in a single serving of popcorn is different from 75, a nutritionist selects a random sample of 20 servings of popcorn and computes the sample mean number of calories to be 78 with a sample standard deviation of 7. Assume the caloric content of popcorn is normally distributed.

I need to compute the test value. I have looked in my book and can't find a similar problem to help me. Any help or show how to get it is greatly appreciated.

Answer by stanbon(75887)

In an article appearing in Today’s Health, a writer states that a single serving of popcorn has on average 75 calories.
To determine if the average number of calories in a single serving of popcorn is different from 75, a nutritionist selects a random sample of 20 servings of popcorn and computes the sample mean number of calories to be 78 with a sample standard deviation of 7. Assume the caloric content of popcorn is normally distributed.
I need to compute the test value.
----
sample mean = x-bar = 78
---
test statistic: t(78) = (78-75)/(7/sqrt(20)) = 1.9166
=============================================
Cheers,
Stan H.